# Fast Revenue Maximization

### v1

@@ 2 Proof of Theorem 1
In this section we prove our main methodological result. Assume $\pazocal{C}$ is equal to $\pazocal{G}$ or $\pazocal{F}$. Fix $N\geq 1$ and let $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ be an information set including $N$ historical prices and conversions rates. Fix a pricing mechanism $\Psi$ in $\pazocal{P}$. We first, note that one can prove the result by considering the more general notion of $\lambda$-regret defined for every $\lambda\in(0,1]$ as,
@@
Taking $\lambda=1$ recovers the absolute regret metric. Furthermore, by using a classical epigraph reformulation (described in Appendix A) of the worst-case ratio one can reduce its analysis to the one of the $\lambda$-regret (see Proposition 1 in Anunrojwong et al. (2023)). In what follows, we prove that the reduction stated in Theorem 1 holds for the $\lambda$-regret.
@@
Notations. We define for every $F$ in $\pazocal{D}_{[\underline{v},\overline{v}]}$, the inverse $F^{-1}(q):=\inf\{v\mbox{ in }[\underline{v},\overline{v}]\mbox{ s.t. }F(v)\geq q\}$ for all $q$ in $[0,1]$. For a distribution with positive density function $f$ on its support $[a,b]$, where $0\leq a<\infty$ and $a\leq b\leq\infty$, we denote the virtual value function for $v\in[a,b]$ by $\phi_{F}(v):=v-\frac{\overline{F}(v)}{f(v)}.$ For any $v\geq 0$, we define $\Gamma(v):=\frac{1}{1+v}$ and $\Gamma^{-1}(q):=\frac{1}{q}-1$ denotes the inverse of $\Gamma$.
@@ 2.1 Candidate worst-case distributions
$\overline{G}_{\pazocal{G}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))$ $=\mathbbm{1}\{v\in[\underline{v},\overline{v}]\}\cdot q_{s^{\prime}}$ (2.1)
$\overline{G}_{\pazocal{F},}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))$ $=\mathbbm{1}\{v\in[\underline{v},\overline{v}]\}\cdot\Gamma\left(\Gamma^{-1}\left(q_{s}\right)+\frac{\Gamma^{-1}\left(q_{s^{\prime}}\right)-\Gamma^{-1}\left(q_{s}\right)}{s^{\prime}-s}(v-s)\right)$ (2.2)
These distributions are motivated by the fact that they allow us to derive the following bounds on distributions in $\pazocal{C}$ by using the so-called single crossing property (Barlow and Proschan, 1975).
@@
Fix $F$ in $\pazocal{C}$ and values $(s,s^{\prime})$ such that $0\leq s\leq s^{\prime}$ and $0<q_{s^{\prime}}=\overline{F}(s^{\prime})\leq q_{s}=\overline{F}(s)$. We have that,
$\overline{F}(v)$ $\geq\overline{G}_{\pazocal{C}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))\qquad\mbox{ if }\>v\mbox{ in }[s,s^{\prime}]$
$\overline{F}(v)$ $\leq\overline{G}_{\pazocal{C}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))\qquad\mbox{ if }\>v>s^{\prime}\text{ or }v<s.$
Lemma 1, proved in Section 5, shows that $\overline{G}_{\pazocal{C}}$ provides a way to upper and lower bound all function in the class $\pazocal{C}$ sharing the same conversion rates at $s$ and $s^{\prime}$. Given that every distribution in $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ have the same conversion rate at prices $\mathbf{p}$, we next use Lemma 1 to construct candidate worst-case distributions. In particular, given a feasible information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i});\quad i=0\cdots N+1\}$ we define for the sake of simple notations, $G_{\pazocal{C},i}(v)={G}_{\pazocal{C}}(v|(p_{i},q_{i}),(p_{i+1},q_{i+1}))$ for every $i\in 0\cdots N$. For both the regular and general classes of distributions, we define on $[\underline{v},\overline{v})$:
$\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})=\overline{G}_{\pazocal{C},i}(v),\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{ for }i=0\cdots N.$ (2.3)
@@
$\overline{U}_{\pazocal{G}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ $=\overline{G}_{\pazocal{G},i-1}(v)=q_{i}\quad\mbox{if }v\mbox{ in}[p_{i},p_{i+1}),\mbox{for }i=0\cdots N.$ (2.4)
$\overline{U}_{\pazocal{F}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ =$=\begin{cases}\min\{1,\overline{G}_{\pazocal{F},1}(v)\}&\quad\mbox{if }v\mbox{ in }[p_{0},p_{1})\\ \min\{\overline{G}_{\pazocal{F},{i-1}}(v),\overline{G}_{\pazocal{F},{i+1}}(v)\}&\quad\mbox{if }v\mbox{ in }[p_{i},p_{i+1}),\mbox{for }i=1\cdots N-2\\ \overline{G}_{\pazocal{F},{i-1}}(v)&\quad\mbox{if }v\mbox{ in }[p_{i},p_{i+1}),\mbox{ for }i\in\{N-1,N\}.\end{cases}$ (2.5)
Intuitively, Lemma 1 implies that the ccdf $\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is a lower bound of the ccdf of any distribution in $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Similarly, the ccdf $\overline{U}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ provides an upper bound. We illustrate this in Figure 1.
@@
Figure 1: Illustration of the implications of the bounds obtained by the single crossing property.
2.2 Reduction to 2D-problem
In this section, we prove our first reduction and develop a reduction of nature’s problem to a two-dimensional optimization problem parameterized by the set of possible values that the optimal price and quantile can take. We first define the set of feasible values given the information at hand. To that end, let ${\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ denote the set of feasible optimal pairs, i.e.,
${\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\left\{({r^{*}},{q^{*}})\mbox{ in }\mathbb{R}_{+}\times[0,1]:\mbox{there exists }\>F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\mbox{ s.t. }{r^{*}}\in\mathscr{O}(F)\mbox{ and }\overline{F}({r^{*}})={q^{*}}\right\}.$
@@ 2.2 Reduction to 2D-problem
$\sup_{F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\left\{\lambda\cdot{r^{*}}\cdot{q^{*}}-\inf_{\begin{subarray}{c}F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):\\ {r^{*}}\in\mathscr{O}(F),\overline{F}({r^{*}})={q^{*}}\end{subarray}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]\right\}$ (2.6)
The above reformulation allows to “decouple” the problem as the first term is fully controlled and second term representing the expected revenue of the pricing mechanism can be optimized independently. Our next structural result solves exactly the inner minimization problem.
@@
$\inf_{\begin{subarray}{c}F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):\\ {r^{*}}\in\mathscr{O}(F),\overline{F}({r^{*}})={q^{*}}\end{subarray}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]=\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)].$
We note that Proposition 1 and (2.6) implies that the problem of nature can be written as
$\sup_{F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathbb{E}_{p\sim\Psi}\left[\lambda\cdot{r^{*}}\cdot{q^{*}}-\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right]$ (2.7)
In other words, given ${r^{*}}\in\mathscr{O}(F)$ and ${q^{*}}$ s.t. ${q^{*}}=\overline{F}({r^{*}})$, one can pin down the worst-case distribution and associated revenue curve.
In the rest of this section, we prove Proposition 1. Our first lemma provides a lower bound on the revenue associated with any value distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
Consider an information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$, then for any pricing mechanism $\Psi\in\pazocal{P}$ and any value distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$, we have that, $\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]\geq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)]$, where $\overline{L}_{\pazocal{C}}$ is defined in (2.3).
@@
Fix $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and consider $v\in[\underline{v},\overline{v}]$. We first show that, $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}).$ Consider $i\in\{0,\ldots,N\}$ such that, $v\in[p_{i},p_{i+1}]$. We remark that $\overline{F}(p_{i})=q_{i}$ and $\overline{F}_{0}(p_{i+1})=q_{i+1}$ because $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Therefore Lemma 1 implies that,
@@
$\mathbb{E}_{\Psi}[Rev\left(p|\overline{F}\right)]=\int_{0}^{\infty}p\cdot\overline{F}(p)\,d\Psi(p)\stackrel{{\scriptstyle(a)}}{{\geq}}\int_{0}^{\infty}p\cdot\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}})\,d\Psi(p)=\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)],$
where $(a)$ holds because $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. This concludes the proof. ∎
@@
First, one can verify that for every $i\in\{0,\ldots,N+1\}$, $\overline{L}_{\pazocal{C}}(p_{i}|\mathscr{I}_{\mathbf{p},\mathbf{q}})=q_{i}.$ Therefore, we need to show that, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}.$
@@
The final structural result needed in the proof of Proposition 1 characterizes the optimal revenue (and associated conversion rate) of $\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
If $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, then $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}.$
@@
We first note that, Lemma 3 implies that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Therefore, for every $i\in\{1,\ldots,N+1\}$, we have $\mathrm{Rev}\left(p_{i}|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=p_{i}\cdot\overline{L}_{\pazocal{C}}(p_{i}|\mathscr{I}_{\mathbf{p},\mathbf{q}})=p_{i}\cdot q_{i}$. Hence, $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}.$
To prove the reverse inequality, it suffices to show that the revenue function $p\mapsto\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)$ is monotonic on every interval $[p_{i},p_{i+1})$ for $i\in\{0,\ldots,N\}$. By the definition in (2.3), we need to analyze the monotonicity of the function $\kappa:p\mapsto p\cdot\overline{G}_{\pazocal{C},i}(p)$.
When $\pazocal{C}=\pazocal{G}$, we have that $\overline{G}_{\pazocal{C},i}(p)$ is constant on $[p_{i},p_{i+1})$. Hence, $\kappa$ is a linear function and is monotonic.
When $\pazocal{C}=\pazocal{F}$, we define for every $q\in[\overline{G}_{\pazocal{C},i}(p_{i}),\overline{G}_{\pazocal{C},i}(p_{i+1})]$ as $\tilde{\kappa}(q)=\overline{G}_{\pazocal{C},i}^{-1}(q)\cdot q$. We note that for every $p\in[p_{i},p_{i+1}]$ we have $\kappa(p)=\tilde{\kappa}(\overline{G}_{\pazocal{C},i}(p))$ and, by composition, if $\tilde{\kappa}$ is monotonic then $\kappa$ is monotonic. To show that $\tilde{\kappa}$ is monotonic we note that it corresponds to the revenue function in the quantile space. Its derivative satisfies, $\frac{d}{dq}\tilde{\kappa}(q)=\phi_{G_{\pazocal{C},i}}(\overline{G}_{\pazocal{C},i}^{-1}(q))$, where $\phi_{F}$ is the virtual value function of a distribution $F$. To conclude the proof, we show in Lemma 6 that the virtual value function of $G_{\pazocal{C},i}$ is constant. Therefore, $\tilde{\kappa}$ is linear and monotonic. ∎
We conclude this section with the proof of Proposition 1.

### v2

@@ 2 Proof of Theorem 1
In this section we prove our main methodological result. Assume $\pazocal{C}$ is equal to $\pazocal{G}$ or $\pazocal{F}$. Fix $N\geq 1$ and let $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ be an information set including $N$ historical prices and conversion rates. Fix a pricing mechanism $\Psi$ in $\pazocal{P}$. We first note that we will prove the result by considering the more general notion of $\lambda$-regret defined for every $\lambda\in[0,1]$ as,
@@
Taking $\lambda=1$ recovers the absolute regret metric. Furthermore, by using a classical epigraph reformulation (described in Section 3) of the worst-case ratio, one can reduce its analysis to the one of the $\lambda$-regret (see Proposition 1 in Anunrojwong et al. (2023)). In what follows, we prove that the reduction stated in Theorem 1 holds for the $\lambda$-regret, for any $\lambda\in[0,1]$.
@@
Notation. We define for every $F$ in $\pazocal{D}_{[\underline{v},\overline{v}]}$, the inverse $F^{-1}(q):=\inf\{v\in[\underline{v},\overline{v}]\mbox{ s.t. }F(v)\geq q\}$ for all $q$ in $[0,1]$. For a distribution with positive density function $f$ on its support $[a,b]$, where $0\leq a<\infty$ and $a\leq b\leq\infty$, we denote the virtual value function for $v\in[a,b]$ by $\phi_{F}(v):=v-\frac{\overline{F}(v)}{f(v)}.$ For any $v\geq 0$, we define $\Gamma(v):=\frac{1}{1+v}$ and $\Gamma^{-1}(q):=\frac{1}{q}-1$ denotes the inverse of $\Gamma$.
@@ 2.1 Candidate worst-case distributions
$\overline{G}_{\pazocal{G}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))$ $=\mathbbm{1}\{v\in[\underline{v},s)\}+\mathbbm{1}\{v\in[s,\overline{v})\}\cdot q_{s^{\prime}}$ (2.1)
$\overline{G}_{\pazocal{F},}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))$ $=\mathbbm{1}\{v\in[\underline{v},\overline{v})\}\cdot\Gamma\left(\Gamma^{-1}\left(q_{s}\right)+\frac{\Gamma^{-1}\left(q_{s^{\prime}}\right)-\Gamma^{-1}\left(q_{s}\right)}{s^{\prime}-s}(v-s)\right).$ (2.2)
These distributions are motivated by the fact that they allow us to derive the following bounds on distributions in $\pazocal{C}$ by deriving a so-called single crossing property.
@@
Fix $F$ in $\pazocal{C}$ and values $(s,s^{\prime})$ such that $\underline{v}\leq s\leq s^{\prime}$ and $0<q_{s^{\prime}}=\overline{F}(s^{\prime}-)\leq q_{s}=\overline{F}(s-)$. We have that,
$\overline{F}(v)$ $\geq\overline{G}_{\pazocal{C}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))\qquad\mbox{ if }\>v\in[s,s^{\prime})$
$\overline{F}(v)$ $\leq\overline{G}_{\pazocal{C}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))\qquad\mbox{ if }\>v\geq s^{\prime}\text{ or }v<s.$
Proof of Lemma 1.
When $\pazocal{C}=\pazocal{F}$, the result follows directly from (Allouah et al., 2022, Lemma 2), which generalizes (Barlow and Proschan, 1975, Chapter 4, Theorem 2.18), with the following additional identity:
When $\pazocal{C}=\pazocal{G}$, the result follows directly from the non-increasing property of complementary cumulative distribution functions. ∎
Lemma 1 shows that $\overline{G}_{\pazocal{C}}$ provides a way to upper and lower bound any function in the class $\pazocal{C}$ sharing the same conversion rates at $s$ and $s^{\prime}$. Given that all distributions in $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ have the same conversion rate at prices $\mathbf{p}$, we next use Lemma 1 to construct candidate worst-case distributions. In particular, given a feasible information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i});\;i\in\{0,\ldots,N+1\}\}$ we define, $G_{\pazocal{C},i}(v)={G}_{\pazocal{C}}(v|(p_{i},q_{i}),(p_{i+1},q_{i+1}))$ for every $i\in\{0,\ldots,N\}$. For both the regular and general classes of distributions, we define on $[\underline{v},\overline{v})$:
$\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})=\overline{G}_{\pazocal{C},i}(v),\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{ for }i\in\{0,\ldots,N\}.$ (2.3)
@@
$\overline{U}_{\pazocal{G}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ $=\overline{G}_{\pazocal{G},i-1}(v)=q_{i}\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{for }i\in\{0,\ldots,N\}.$
$\overline{U}_{\pazocal{F}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ =$=\begin{cases}\min\{1,\overline{G}_{\pazocal{F},1}(v)\}&\quad\mbox{if }v\in[p_{0},p_{1})\\ \min\{\overline{G}_{\pazocal{F},{i-1}}(v),\overline{G}_{\pazocal{F},{i+1}}(v)\}&\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{for }i\in\{1,\ldots,N-2\}\\ \overline{G}_{\pazocal{F},{i-1}}(v)&\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{ for }i\in\{N-1,N\}.\end{cases}$
Intuitively, Lemma 1 implies that the ccdf $\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is a lower bound of the ccdf of any distribution in $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Similarly, the ccdf $\overline{U}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ provides an upper bound. We illustrate this in Figure 1. The information set is $\{(p_{i},q_{i}),\,i\in\{1,\ldots,4\}\}=\{(0.47,0.25),(0.70,0.055),(0.89,0.017),(0.98,0.006)\}$ and we set $\underline{v}=0,\,\overline{v}=1$.
@@
Figure 1: Illustration of the implications of the bounds derived in Lemma 1.
Figure 1-a depicts in red the information set. The shaded region in green (resp. in blue) depicts the region where the ccdf of regular (resp. general) distributions that are consistent with the information set could fall. Figure 1-b depicts the corresponding region for the revenue curve.
2.2 Reduction to a two-dimensional problem
In this section, we prove our first reduction of Nature’s problem to a two-dimensional optimization problem parameterized by the set of possible values that the optimal price and corresponding conversion rate can take. To that end, let ${\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ denote the set of feasible optimal pairs, i.e.,
${\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\left\{({r^{*}},{q^{*}})\in[\underline{v},\overline{v}]\times[0,1]:\mbox{there exists }\>F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\mbox{ s.t. }{r^{*}}\in\mathscr{O}(F)\mbox{ and }\overline{F}({r^{*}}-)={q^{*}}\right\}.$
@@ 2.2 Reduction to a two-dimensional problem
$\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\left\{\lambda\cdot{r^{*}}\cdot{q^{*}}-\inf_{\begin{subarray}{c}F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):\\ {r^{*}}\in\mathscr{O}(F),\overline{F}({r^{*}}-)={q^{*}}\end{subarray}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]\right\}$ (2.4)
The above reformulation allows to “decouple” the problem as the first term is fully controlled by ${r^{*}}$ and ${q^{*}}$. The second term, which represents the expected revenue of the pricing mechanism, can then be optimized independently. Our next structural result solves exactly the inner minimization problem.
@@
$\inf_{\begin{subarray}{c}F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):\\ {r^{*}}\in\mathscr{O}(F),\overline{F}({r^{*}}-)={q^{*}}\end{subarray}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]=\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)].$
We note that Proposition 1 and (2.4) imply that Nature’s problem can be written as
$\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathbb{E}_{p\sim\Psi}\left[\lambda\cdot{r^{*}}\cdot{q^{*}}-\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right].$ (2.5)
In other words, given ${r^{*}}\in\mathscr{O}(F)$ and ${q^{*}}$ s.t. ${q^{*}}=\overline{F}({r^{*}}-)$, one can pin down the worst-case distribution and associated revenue curve. In what follows, we prove Proposition 1.
Our first lemma shows that our candidate worst-case distribution provides a lower bound on the revenue associated with any value distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
Consider an information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$, then for any pricing mechanism $\Psi\in\pazocal{P}$ and any value distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$, we have that, $\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]\geq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)]$, where the ccdf of $L_{\pazocal{C}}$ is defined in (2.3).
@@
Fix $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and consider $v\in[\underline{v},\overline{v}]$. We first show that, $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}).$ Consider $i\in\{0,\ldots,N\}$ such that, $v\in[p_{i},p_{i+1})$. We remark that $\overline{F}(p_{i}-)=q_{i}$ and $\overline{F}_{0}(p_{i+1}-)=q_{i+1}$ because $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Therefore Lemma 1 implies that,
@@
$\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]=\int_{0}^{\infty}p\cdot\overline{F}(p-)\,d\Psi(p)\stackrel{{\scriptstyle(a)}}{{\geq}}\int_{0}^{\infty}p\cdot\overline{L}_{\pazocal{C}}(p-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\,d\Psi(p)=\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)],$
where $(a)$ holds because $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ for every $v\in[\underline{v},\overline{v}]$ and therefore by taking a right limit $\overline{F}(v-)\geq\overline{L}_{\pazocal{C}}(v-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ for every $v\in[\underline{v},\overline{v}]$. This concludes the proof. ∎
@@
First, one can verify that for every $i\in\{0,\ldots,N+1\}$, $\overline{L}_{\pazocal{C}}(p_{i}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})=q_{i}.$ Therefore, we need to show that, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}.$
@@
The final structural result needed in the proof of Proposition 1 characterizes the optimal revenue (and associated conversion rate) for the distribution $\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
If $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, then $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}.$
@@
We first note that Lemma 3 implies that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Therefore, for every $i\in\{1,\ldots,N+1\}$, we have $\mathrm{Rev}\left(p_{i}|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=p_{i}\cdot\overline{L}_{\pazocal{C}}(p_{i}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})=p_{i}\cdot q_{i}$. Hence, $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}.$
To prove the reverse inequality, it suffices to show that the revenue function $p\mapsto\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)$ is monotonic on every interval $[p_{i},p_{i+1})$ for $i\in\{0,\ldots,N\}$. By the definition in (2.3), we need to analyze the monotonicity of the function $\kappa:p\mapsto p\cdot\overline{G}_{\pazocal{C},i}(p)$.
When $\pazocal{C}=\pazocal{G}$, we have that $\overline{G}_{\pazocal{C},i}(p)$ is constant on $[p_{i},p_{i+1})$. Hence, $\kappa$ is a linear function and is monotonic on $[p_{i},p_{i+1})$.
When $\pazocal{C}=\pazocal{F}$, we define for every $q\in[\overline{G}_{\pazocal{C},i}(p_{i}),\overline{G}_{\pazocal{C},i}(p_{i+1}))$ the function $\tilde{\kappa}(q)=\overline{G}_{\pazocal{C},i}^{-1}(q)\cdot q$. We note that for every $p\in[p_{i},p_{i+1})$ we have $\kappa(p)=\tilde{\kappa}(\overline{G}_{\pazocal{C},i}(p))$ and, by composition, if $\tilde{\kappa}$ is monotonic then $\kappa$ is monotonic. To show that $\tilde{\kappa}$ is monotonic we note that it corresponds to the revenue function in the quantile space. Its derivative satisfies, $\frac{d}{dq}\tilde{\kappa}(q)=\phi_{G_{\pazocal{C},i}}(\overline{G}_{\pazocal{C},i}^{-1}(q))$, where $\phi_{F}$ is the virtual value function of a distribution $F$. To conclude the proof, we show in Lemma 6 that the virtual value function of $G_{\pazocal{C},i}$ is constant. Therefore, $\tilde{\kappa}$ is linear and monotonic on $[q_{i},q_{i+1})$. ∎
We next combine these lemmas to formally prove Proposition 1.